A delivery truck hits a parked car from behind. The mass of the car is 2550 kg. What force is exerted on the car if (a) the car is rigid and accelerates to 4.5 m/s in 0.20 s? (b) the car has crumple zones and accelerates to 4.5 m/s in 0.30 s?
1
Expert's answer
2020-03-25T11:07:55-0400

a)


"F=\frac{mv}{t}=2550\frac{4.5}{0.2}=57000\ N=57\ kN"

b)


"F=\frac{mv}{t'}=2550\frac{4.5}{0.3}=38000\ N=38\ kN"